# Enhancing Speech Large Language Models with Prompt-Aware Mixture of Audio Encoders

### v2

@@ Abstract
Connecting audio encoders with large language models (LLMs) allows the LLM to perform various audio understanding tasks, such as automatic speech recognition (ASR) and audio captioning (AC). Most research focuses on training an adapter layer to generate a unified audio feature for the LLM. However, different tasks may require distinct features that emphasize either semantic or acoustic aspects, making task-specific audio features more desirable. In this paper, we propose Prompt-aware Mixture (PaM) to enhance the Speech LLM that uses multiple audio encoders. Our approach involves using different experts to extract different features based on the prompt that indicates different tasks. Experiments demonstrate that with PaM, only one Speech LLM surpasses the best performances achieved by all single-encoder Speech LLMs on ASR, Speaker Number Verification, and AC tasks. PaM also outperforms other feature fusion baselines, such as concatenation and averaging. Our code would be available at: https://github.com/shanweiqiao/PaM
@@ 1 Introduction
In this paper, we introduce Prompt-aware Mixture (PaM), a novel MoE method for merging multiple encoders to enhance Speech LLMs. Our approach integrates a prompt-aware routing mechanism, emphasizes feature fusion, and considers the relative importance of each encoder for different tasks, aiming to improve all downstream performance. PaM employs three distinct audio encoders: the Whisper encoder, WavLM, and Wav2Vec2 Baevski et al. (2020). We train a set of experts for prompt-aware feature fusion, comprising one shared expert and four task-specific experts. On each task, an expert learns the optimal weights for each encoder and its respective layers, and subsequently maps the resulting features to the embedding space of the Qwen2.5 model Team (2024). The embedding of the prompt is utilized to determine the appropriate routing. Notably, in PaM, the routing guides the selection of the fusion parameters rather than the choice of the encoder. Experiments are conducted across three tasks: ASR, speaker number verification(SNV), and AC. On all datasets, including LibriSpeech Panayotov et al. (2015), AMI Kraaij et al. (2005), AIR-Bench (SNV) Yang et al. (2024), and AudioCaps Kim et al. (2019), PaM achieves relative improvements of 15%, 25%, 3.4%, and 7.6%, respectively, in comparison to the best-performing single-encoder Speech LLM, and achieves a better average rank on all tasks compared with conventional approaches. Our contributions can be summarized as follows:
@@
  • We introduce PaM, a specialized MoE method that incorporates a prompt-aware routing mechanism to assign distinct weights to each encoder and its layers based on the task.
@@
Figure 2: The architecture of the proposed PaM method. The output feature $\mathbf{H}_{\text{prompt}}$ of the prompt $\mathbf{X}_{\text{prompt}}$ guides the routing of the MoE adapter, which incorporates a fixed shared expert (denoted as expert S) and a single routed expert (denoted as expert expert $i\in[1,N]$, where NN represents the total number of routed experts). For each expert, the last hidden states from all encoders $\mathbf{h}_{i}^{L_{i}-1}$ are concatenated with $K$ fused hidden states ($K$ being a hyperparameter, with the default value $K=3$) derived from a fusion weight matrix $\mathbf{W}$. Subsequently, a feedforward network (FFN) is applied to align with the dimensions of the LLM.
@@ 2 Method
In this section, we begin with an overview of the proposed PaM method (Figure 2). We then elaborate on the details of the encoder fusion process, executed by a single expert, and describe the prompt-aware routing method.
@@ 2.1 Overall Architecture
Next, as shown in Equation 2.1, we combine these features using an MoE fusion method, which includes a shared expert and N routed experts where the number of routed experts corresponds to the number of predefined tasks. During inference, only one routed expert is selected, based on the task indicated by the prompt.
@@ 2.2 Multi-layer Fusion
$\mathbf{H}_{i}=\{\mathbf{h}_{i}^{0},\mathbf{h}_{i}^{1},...\mathbf{h}_{i}^{L_{i}-1}\}$ (4)
As illustrated in Equation 5, we consider the hidden states $\mathbf{h}_{i}^{\{0,1,\ldots,(L_{i}\text{-}2)\}}$ to derive the fused hidden states $\mathbf{h}_{k}^{\text{fused}}$. For each hidden state $\mathbf{h}^{l}_{i}$ of all three encoders, a set of scalar weights containing $\sum_{i}^{3}(L_{i}-1)$ elements is assigned to control its relative importance. We denote a set of scalar weights as $\{w_{i,k}^{l}|i\in[1,3],l\in[1,L_{i}-1]\}$ and then used it to generate fused hidden states $\mathbf{h}_{k}^{\text{fused}}$. To maintain the diversity of the fusion feature, we utilize $K$ sets of scalar weights. Thus, the dimension of the whole learnable matrices is $\mathbf{W}\in\mathbb{R}^{K\times(\sum_{i}^{3}(L_{i}-1))}$.
@@ 2.3 Prompt Aware Routing
A prompt refers to a text segment that provides context or objectives for generation, which can typically be categorized into several distinct types according to task. For instance, speech-related tasks, sound-related tasks, and speech chat tasks Yang et al. (2024). In this paper, we investigate three tasks: ASR, speaker number verification, and AC. We utilize distinct experts for each task. For effective routing, the router must identify the task type based on the prompt. We employ a simple classification approach (Equation 7 and 8) wherein we use the last hidden states $\mathbf{H}_{\text{prompt}}$ of the prompt from the LLM, followed by a FFN and Softmax activation, to obtain the task posteriors $\text{P}(\text{Task}|\mathbf{H}_{\text{prompt}})$.
@@ 2.4 Training Objective
The training loss function, as illustrated in Equation 2.4, is the sum of the cross-entropy loss $\mathcal{L}_{\text{G}}$ for prompt-aware routing and the cross-entropy loss $\mathcal{L}_{\text{llm}}$ between the LLM’s output $\mathbf{Y}$ and the /ground truth $\hat{\mathbf{Y}}$. Loss function, as illustrated in Equation 2.4, is the sum of the cross-entropy loss $\mathcal{L}_{\text{G}}$ for prompt-aware routing and the cross-entropy loss $\mathcal{L}_{\text{llm}}$ between the LLM’s output $\mathbf{Y}$ and the ground truth $\hat{\mathbf{Y}}$.
@@ 3.1 Datasets and Evaluation Metrics
We assess the efficacy of our method across three audio-to-text tasks: automatic speech recognition (ASR), speaker number verification (SNV), and audio captioning (AC). We list the detailed information of training data in Appendix A.2. In total, the training data contains 450 hours of audio signals. The test dataset includes LibriSpeech-test-clean, LibriSpeech-test-other, AMI, the SNV test set from AIR-Bench, and the test set of AudioCaps along with its corresponding QA version from AudioBench Wang et al. (2024a), which contains diverse questions. ASR tasks focus on semantics. The LibriSpeech test set originates from audiobooks, demonstrating ASR performance in a clean scenario. AMI, a real meeting corpus containing spontaneous talk, reflects ASR performance in a more challenging, real-world scenario. SNV and AC test sets can indicate the Speech LLM’s ability to understand speaker and acoustic information. We evaluate the performance using word error rate (WER) for ASR tasks, accuracy for SNV, and METEOR Banerjee and Lavie (2005) for AC. Additionally, we list the results on AC and AC QA tasks with more metrics in Appendix A.4.
@@
   - Whisper Radford et al. (2022) 9.61 16.73 16.27 18.80% 32.96 15.04 5.67
   - WavLM Chen et al. (2022) 5.59 10.57 18.97 41.40% 27.14 12.77 5.50
   - Wav2Vec2 Baevski et al. (2020) 4.30 9.46 26.69 39.00% 23.81 11.20 5.33
@@
   - WavLLM Hu et al. (2024) 4.95 (- 0.65) 9.19 (+0.27) 15.29 (+0.98) 39.20% (- 2.20) 34.93 (+1.97) 16.35 (+1.31) 2.67
   - SALMONN Tang et al. (2024) 5.04 (- 0.74) 9.70 (- 0.24) 19.04 (- 2.77) 49.40% (+8.00) 34.86 (+1.90) 15.97 (+0.93) 3.50
   - Average 4.76 (- 0.46) 10.43 (- 0.97) 17.20 (- 0.93) 45.50% (+4.10) 33.22 (+0.26) 15.53 (+0.49) 3.67
PaM (Ours) 3.65 (+0.65) 7.07 (+2.39) 12.79 (+3.48) 42.80% (+1.40) 35.47 (+2.51) 15.70 (+0.66) 1.67
@@ 3.2 Model Architecture and Training
We train our model based on Huggingface Transformers Library [^4] https://github.com/huggingface/transformers. Our model consists of three audio encoders, a pre-fusion adapter for each encoder, a PaM fusion module, and an LLM. In our main experiments, the encoders are Whisper-Small encoder, WavLM-Base-Plus, and Wav2Vec2-Base-960h [^5] The links to the pretrained models and datasets used can be found in Appendix A.2. The implementation details of baseline methods are available in Appendix A.3., each with approximately 100 million parameters. We downsample the features from the Whisper encoder by a factor of two, resulting in a frame length of 40ms, consistent with the frame length of Wav2Vec2 and WavLM. The pre-fusion adapter is an FFN that transforms the encoder’s hidden dimension $D_{\text{E}}$ to the LLM’s hidden dimension $D_{\text{LLM}}$. Each expert in the PaM fusion module includes a fusion weight matrix ($\mathbb{R}^{3L\times 3}$) and a linear layer ($\mathbb{R}^{6D_{\text{LLM}}\times D_{\text{LLM}}}$) to fuse features from all encoders. Here, $L$ represents the number of layers in the encoder, which is 12 for all encoders in our experiments. For the fused features, we set $K=3$, corresponding to the number of last hidden states. We utilize four routed experts, each corresponding to a specific task category: ASR-clean, ASR-noisy, SNV, and AC. For each category, we generate 50 prompts using ChatGPT OpenAI (2023) [^6] We provide few examples of the prompts we used in Appendix A.11. For the LLM model, we select the Qwen2.5-3B Team (2024). In section 4, we also experiment with other encoders, including Hubert-Base-LS960, Whisper-Large-v3, and WavLM-Large. [^7] Additionally, we add the BEATs model, which performs well on the AC task, to further enhance PaM in Appendix A.5.
@@ 4.1 Main Results
PaM consistently outperforms all single encoder baselines, delivering performance improvements across all tasks. This consistent improvement can be attributed to the MoE adapter, which provides unique features tailored for each task. Compared to other fusion methods (i.e., concatenation and averaging), PaM achieves significantly lower WERs on the LibriSpeech and AMI datasets and similar performance on SNV and AC tasks.
@@ 4.2 Feature Importance
In Figure 3, we visualize the fusion weights for each expert, excluding the shared expert, which can be interpreted as the fusion weight for each task. We summed the weights for every four layers to enhance clarity, resulting in the total weight for shallow, middle, and deep layers. Generally, different tasks require different features, so each expert has distinct fusion weights. Specifically, when the expert for ASR-clean is activated, it mainly focuses on features from WavLM and Wav2Vec2, especially the deep layers. When the expert for ASR-noise is activated, it primarily focuses on features from the Whisper encoder and WavLM. For SNV and AC tasks, all three encoders have similar fusion weights. For the SNV task, features from the middle layers are more important, while for the AC task, shallow layers contribute more.
@@ 4.3 Combinations of Encoders
We can observe that different encoders offer varying benefits, which proves that the task-specific encoders can significantly improve the performance on the corresponding tasks. For example, when Whisper is used, regardless of how many encoders are employed, the Speech LLM achieves the lowest WER on AMI and the highest METEOR scores on AudioCaps. On the other hand, Wav2Vec2 provides an advantage for recognizing speech signals in LibriSpeech. This indicates that when selecting encoders for Speech LLM, it is essential to consider the domain, downstream tasks, and the capabilities of each encoder. It is suggested to use a robust general domain model like Whisper in combination with domain-specific encoders such as Wav2Vec2.
@@ 4.4 Larger Encoders and LLMs
We try to further enhance performance by replacing the encoders in the proposed method with their larger versions (Table 3). Specifically, we replace the Whisper-Small encoder with the Whisper-Medium encoder, Wav2Vec2-Base-960h with Wav2Vec2-Large-960h, and WavLM-Base-Plus with WavLM-Large. Our observations indicate that on the LibriSpeech-clean dataset, performance does not significantly improve and may even slightly degrade. However, for the SNV and AC tasks, performance consistently improves, suggesting that more challenging sound-related tasks benefit more from better encoders. Additionally, we observe that when all encoders are replaced with their larger versions, we achieve the best performance across almost all tasks, albeit at the cost of increased computation.
In our investigation of other LLMs, including Qwen2.5-7B, LLaMA3.2-3B, and LLaMA3.1-8B, we observed some improvements in certain tasks. However, the overall performance was not superior to that of Qwen2.5-3B. The potential reason for this is that we used short audios, and both the prompts and answers were brief, thereby not fully utilizing the strong semantic understanding capabilities of the larger LLMs. Consequently, we opted to use Qwen2.5-3B in this paper. It is important to emphasize that for Speech LLMs, the extracted features may be more critical than the LLM itself for many downstream tasks. In addition, we also compare the concatenation fusion strategy (WavLLM) and PaM using larger LLMs based on the LLaMA3.1-8B and Qwen2.5-7B. We find that the performance of the concatenation fusion strategy is inconsistent across different base models of similar size, whereas PaM maintains stability, as detailed in Appendix A.7.

### v3

@@ Abstract
Connecting audio encoders with large language models (LLMs) allows the LLM to perform various audio understanding tasks, such as automatic speech recognition (ASR) and audio captioning (AC). Most research focuses on training an adapter layer to generate a unified audio feature for the LLM. However, different tasks may require distinct features that emphasize either semantic or acoustic aspects, making task-specific audio features more desirable. In this paper, we propose Prompt-aware Mixture (PaM) to enhance the Speech LLM that uses multiple audio encoders. Our approach involves using different experts to extract different features based on the prompt that indicates different tasks. Experiments demonstrate that with PaM, only one Speech LLM surpasses the best performances achieved by all single-encoder Speech LLMs on ASR, Speaker Number Verification, and AC tasks. PaM also outperforms other feature fusion baselines, such as concatenation and averaging. Our code will be available at: https://github.com/shanweiqiao/PaM
@@ 1 Introduction
In this paper, we introduce Prompt-aware Mixture (PaM), a novel method based on MoE for merging multiple encoders to enhance Speech LLMs. Our approach integrates a prompt-aware routing mechanism, emphasizes feature fusion, and considers the relative importance of each encoder for different tasks, aiming to improve all downstream performance. PaM employs three distinct audio encoders: the Whisper encoder, WavLM, and Wav2Vec2 Baevski et al. (2020). We train a set of experts for prompt-aware feature fusion, comprising one shared expert and four task-specific experts. On each task, an expert learns the optimal weights for each encoder and its respective layers, and subsequently maps the resulting features to the embedding space of the Qwen2.5 model Team (2024). The embedding of the prompt is utilized to determine the appropriate routing. Notably, in PaM, the routing guides the selection of the fusion parameters rather than the choice of the encoder. Experiments are conducted across three tasks: ASR, speaker number verification (SNV), and AC. On all datasets, including LibriSpeech Panayotov et al. (2015), AMI Kraaij et al. (2005), AIR-Bench (SNV) Yang et al. (2024), and AudioCaps Kim et al. (2019), PaM achieves relative improvements of 15%, 25%, 3.4%, and 7.6%, respectively, in comparison to the best-performing single-encoder Speech LLM, and achieves a better average rank on all tasks compared with conventional approaches. Our contributions can be summarized as follows:
@@
  • We introduce PaM, a method based on MoE that incorporates a prompt-aware routing mechanism to assign distinct weights to each encoder and its layers based on the task.
@@
Figure 2: The architecture of the proposed PaM method. The output feature $\mathbf{H}_{\text{prompt}}$ of the prompt $\mathbf{X}_{\text{prompt}}$ guides the routing of the MoE structure adapter, which incorporates a fixed shared expert (denoted as expert S) and a single routed expert (denoted as expert $i\in[1,N]$, where NN represents the total number of routed experts). For each expert, the last hidden states from all encoders $\mathbf{h}_{i}^{L_{i}-1}$ are concatenated with $K$ fused hidden states ($K$ being a hyperparameter, with the default value $K=3$) derived from a fusion weight matrix $\mathbf{W}$. Subsequently, a feedforward network (FFN) is applied to align with the dimensions of the LLM.
@@ 2 Method
In this section, we begin with an overview of the proposed PaM method (Figure 2). We then elaborate on the details of the encoder fusion process executed by a single expert, and describe the prompt-aware routing method.
@@ 2.1 Overall Architecture
Next, as shown in Equation 2.1, we combine these features using a fusion method based on MoE, which includes a shared expert and N routed experts, where the number of routed experts corresponds to the number of predefined tasks. During inference, only one routed expert is selected, based on the task indicated by the prompt.
@@ 2.2 Multi-layer Fusion
$\mathbf{H}_{i}=\{\mathbf{h}_{i}^{0},\mathbf{h}_{i}^{1},\ldots\mathbf{h}_{i}^{L_{i}-1}\}$ (4)
As illustrated in Equation 5, we consider the hidden states $\mathbf{h}_{i}^{\{0,1,\ldots,(L_{i}\text{-}2)\}}$ to derive the fused hidden states $\mathbf{h}_{k}^{\text{fused}}$. For each hidden state $\mathbf{h}^{l}_{i}$ of all three encoders, a set of scalar weights containing $\sum_{i}^{3}(L_{i}-1)$ elements is assigned to control its relative importance. We denote a set of scalar weights as $\{w_{i,k}^{l}|i\in[1,3],l\in[1,L_{i}-1]\}$ and then use it to generate fused hidden states $\mathbf{h}_{k}^{\text{fused}}$. To maintain the diversity of the fusion feature, we utilize $K$ sets of scalar weights. Thus, the dimension of the whole learnable matrices is $\mathbf{W}\in\mathbb{R}^{K\times(\sum_{i}^{3}(L_{i}-1))}$.
@@ 2.3 Prompt Aware Routing
A prompt refers to a text segment that provides context or objectives for generation, which can typically be categorized into several distinct types according to task. For instance, speech-related tasks, sound-related tasks, and speech chat tasks Yang et al. (2024). In this paper, we investigate three tasks: ASR, SNV, and AC. We utilize distinct experts for each task. For effective routing, the router must identify the task type based on the prompt. We employ a simple classification approach (Equation 7 and 8) wherein we use the last hidden states $\mathbf{H}_{\text{prompt}}$ of the prompt from the LLM, followed by an FFN and Softmax activation, to obtain the task posteriors $\text{P}(\text{Task}|\mathbf{H}_{\text{prompt}})$.
@@ 2.4 Training Objective
The training loss function, as illustrated in Equation 2.4, is the sum of the cross-entropy loss $\mathcal{L}_{\text{G}}$ for prompt-aware routing and the cross-entropy loss $\mathcal{L}_{\text{llm}}$ between the LLM’s output $\mathbf{Y}$ and the ground truth $\hat{\mathbf{Y}}$. $\hat{\mathbf{Y}}$.
@@ 3.1 Datasets and Evaluation Metrics
We assess the efficacy of our method across three audio-to-text tasks: automatic speech recognition (ASR), speaker number verification (SNV), and audio captioning (AC). We list the detailed information on training data in Appendix A.2. In total, the training data contains 450 hours of audio signals. The test dataset includes LibriSpeech-test-clean, LibriSpeech-test-other, AMI, the SNV test set from AIR-Bench, and the test set of AudioCaps along with its corresponding QA version from AudioBench Wang et al. (2024a), which contains diverse questions. ASR tasks focus on semantics. The LibriSpeech test set originates from audiobooks, demonstrating ASR performance in a clean scenario. AMI, a real meeting corpus containing spontaneous talk, reflects ASR performance in a more challenging, real-world scenario. SNV and AC test sets can indicate the Speech LLM’s ability to understand speaker and acoustic information. We evaluate the performance using word error rate (WER) for ASR tasks, accuracy for SNV, and METEOR Banerjee and Lavie (2005) for AC. Additionally, we list the results on AC and AC QA tasks with more metrics in Appendix A.4.
@@
   - Whisper Radford et al. (2022) 9.61 16.73 16.27 18.8% 32.96 15.04 5.67
   - WavLM Chen et al. (2022) 5.59 10.57 18.97 41.4% 27.14 12.77 5.50
   - Wav2Vec2 Baevski et al. (2020) 4.30 9.46 26.69 39.0% 23.81 11.20 5.33
@@
   - WavLLM Hu et al. (2024) 4.95 (- 0.65) 9.19 (+0.27) 15.29 (+0.98) 39.2% (- 2.20) 34.93 (+1.97) 16.35 (+1.31) 2.67
   - SALMONN Tang et al. (2024) 5.04 (- 0.74) 9.70 (- 0.24) 19.04 (- 2.77) 49.4% (+8.00) 34.86 (+1.90) 15.97 (+0.93) 3.50
   - Average 4.76 (- 0.46) 10.43 (- 0.97) 17.20 (- 0.93) 45.5% (+4.10) 33.22 (+0.26) 15.53 (+0.49) 3.67
PaM (Ours) 3.65 (+0.65) 7.07 (+2.39) 12.79 (+3.48) 42.8% (+1.40) 35.47 (+2.51) 15.70 (+0.66) 1.67
@@ 3.2 Model Architecture and Training
We train our model based on Huggingface Transformers Library [^4] https://github.com/huggingface/transformers. Our model consists of three audio encoders, a pre-fusion adapter for each encoder, a PaM fusion module, and an LLM. In our main experiments, the encoders are Whisper-Small encoder, WavLM-Base-Plus, and Wav2Vec2-Base-960h [^5] The links to the pretrained models and datasets used can be found in Appendix A.2. The implementation details of baseline methods are available in Appendix A.3., each with approximately 100 million parameters. We downsample the features from the Whisper encoder by a factor of two, resulting in a frame length of 40ms, consistent with the frame length of Wav2Vec2 and WavLM. The pre-fusion adapter is an FFN that transforms the encoder’s hidden dimension $D_{\text{E}}$ to the LLM’s hidden dimension $D_{\text{LLM}}$. Each expert in the PaM fusion module includes a fusion weight matrix ($\mathbb{R}^{3L\times 3}$) and a linear layer ($\mathbb{R}^{6D_{\text{LLM}}\times D_{\text{LLM}}}$) to fuse features from all encoders. Here, $L$ represents the number of layers in the encoder, which is 12 for all encoders in our experiments. For the fused features, we set $K=3$, corresponding to the number of last hidden states. We utilize four routed experts, each corresponding to a specific task category: ASR-clean, ASR-noisy, SNV, and AC. For each category, we generate 50 prompts using ChatGPT OpenAI (2023) [^6] We provide a few examples of the prompts we used in Appendix A.11. For the LLM model, we select the Qwen2.5-3B Team (2024). In Section 4, we also experiment with other encoders, including Hubert-Base-LS960, Whisper-Large-v3, and WavLM-Large. [^7] Additionally, we add the BEATs model, which performs well on the AC task, to further enhance PaM in Appendix A.5.
@@ 4.1 Main Results
PaM consistently outperforms all single encoder baselines, delivering performance improvements across all tasks. This consistent improvement can be attributed to the MoE structure adapter, which provides unique features tailored for each task. Compared to other fusion methods (i.e., concatenation and averaging), PaM achieves significantly lower WERs on the LibriSpeech and AMI datasets and similar performance on SNV and AC tasks.
@@ 4.2 Feature Importance
In Figure 3, we visualize the fusion weights for each expert, excluding the shared expert, which can be interpreted as the fusion weight for each task. We sum the weights for every four layers to enhance clarity, resulting in the total weight for shallow, middle, and deep layers. Generally, different tasks require different features, so each expert has distinct fusion weights. Specifically, when the expert for ASR-clean is activated, it mainly focuses on features from WavLM and Wav2Vec2, especially the deep layers. When the expert for ASR-noise is activated, it primarily focuses on features from the Whisper encoder and WavLM. For SNV and AC tasks, all three encoders have similar fusion weights. For the SNV task, features from the middle layers are more important, while for the AC task, shallow layers contribute more.
@@ 4.3 Combinations of Encoders
We can observe that different encoders offer varying benefits, which proves that the task-specific encoders can significantly improve the performance on the corresponding tasks. For example, when Whisper is used, regardless of how many encoders are employed, the Speech LLM achieves the lowest WER on AMI and the highest METEOR scores on AudioCaps. On the other hand, Wav2Vec2 provides an advantage for recognizing speech signals in LibriSpeech. This indicates that when selecting encoders for Speech LLM, it is essential to consider the domain, downstream tasks, and the capabilities of each encoder. We suggest using a robust general domain model like Whisper in combination with domain-specific encoders such as Wav2Vec2.
@@ 4.4 Larger Encoders and LLMs
We aim to further enhance performance by replacing the encoders in the proposed method with their larger versions (Table 3). Specifically, we replace the Whisper-Small encoder with the Whisper-Medium encoder, Wav2Vec2-Base-960h with Wav2Vec2-Large-960h, and WavLM-Base-Plus with WavLM-Large. Our observations indicate that on the LibriSpeech-clean dataset, performance does not significantly improve and may even slightly degrade. However, for the SNV and AC tasks, performance consistently improves, suggesting that more challenging sound-related tasks benefit more from better encoders. Additionally, we observe that when all encoders are replaced with their larger versions, we achieve the best performance across almost all tasks, albeit at the cost of increased computation.
In our investigation of other LLMs, including Qwen2.5-7B, LLaMA3.2-3B, and LLaMA3.1-8B, we observed some improvements in certain tasks. However, the overall performance was not superior to that of Qwen2.5-3B. The potential reason for this is that we used short audios samples, and both the prompts and answers were brief, thereby not fully utilizing the strong semantic understanding capabilities of the larger LLMs. Consequently, we opted to use Qwen2.5-3B in this paper. It is important to emphasize that for Speech LLMs, the extracted features may be more critical than the LLM itself for many downstream tasks. In addition, we also compare the concatenation fusion strategy (WavLLM) and PaM using larger LLMs based on the LLaMA3.1-8B and Qwen2.5-7B. We find that the performance of the concatenation fusion strategy is inconsistent across different base models of similar size, whereas PaM maintains stability, as detailed in Appendix A.7.